OpenStudy (anonymous):

I need help as to where I went wrong please! I will post the question as a comment where I can use the equation thing.

OpenStudy (anonymous):

10. (i) Find the first three terms, in descending powers of x, in the expansion of \[( x - \frac{ 2 }{ x })^5\] (ii) Hence find the coefficient of x in the expansion of \[(4 + \frac{ 1 }{ x^2 })(x - \frac{ 2 }{ x })^5\]

OpenStudy (anonymous):

I need help with (ii), I got the answers to the first part correct

OpenStudy (anonymous):

I multiplied the three terms from (i) by \[(4 + \frac{ 1 }{ x^2 })\]

OpenStudy (anonymous):

And I got 160 but the correct answer is 150, can you tell me what I did wrong?

OpenStudy (whpalmer4):

First 3 terms in the expansion of (i) are \[x^5-10x^3+40x\]right? Multiply that by \[4+\frac{1}{x^2} = 4+x^{-2}\] \[(4+x^{-2})(x^5-10x^3+40x) = 4(x^5-10x^3+40x)+x^{-2}(x^5-10x^3+40x)\]ignoring the terms that don't end up as \(x\) terms: \[4(40x)+x^{-2}(-10x^3) = 160x-10x^1=160x-10x = 150x\] My guess is that you made a mistake when you multiplied \(\frac{1}{x^2}*(-10x^3)\) or didn't collect like terms...

OpenStudy (anonymous):

Yes I made a mistake with multiplication I think, thank you very much!